# Learning Contrastive Feature Representations for Facial Action Unit Detection

### v5

@@ 4.2.4 Evaluation on Aff-Wild2
AUNCE $\textbf{64.1}^{{\color[rgb]{1,0,0}1.7\uparrow}}$ $\textbf{80.2}^{{\color[rgb]{1,0,0}1.6\uparrow}}$
MEGraph [33] 2022IJCAI WCE 63.6 80.0
(Onestage) AUNCE $\textbf{66.1}^{{\color[rgb]{1,0,0}2.5\uparrow}}$ $\textbf{80.4}^{{\color[rgb]{1,0,0}0.4\uparrow}}$
@@ 4.3 Ablation Study
4.3.1 Compared with weighted cross-entropy loss
Due to the primary innovation of this paper being the introduction of a novel contrastive loss, AUNCE, which is compatible with various AU detection encoders, we conducted a comparative analysis between the proposed AUNCE loss and the traditional weighted cross-entropy loss (WCE) across multiple AU detection encoders, as detailed in Table 6. The findings reveal that AUNCE demonstrates superior performance in AU detection. Notably, the MEGraph encoder achieved the best results, with the F1-score on the BP4D dataset improving from 63.6% to 66.1%. This underscores the validity of capturing unique features of different samples over focusing solely on pixel-level information of the entire face associated with AU.
Table 7: Comparison of the efficiency of WCE loss and AUNCE loss on BP4D and DISFA dataset, where encoders of these two losses are both MEGraph with the first-stage training network.
Dataset Loss Frame Rate(FPS) 1 Epoch Time(s)
BP4D WCE 0.0271 2734
AUNCE 0.0138 1388
DISFA WCE 0.0265 2308
AUNCE 0.0136 1189
Furthermore, as shown in Table 7, when the number of encoder parameters is the same, our proposed AUNCE loss increases the frame rate and reduces the training time of one epoch by nearly half compared to WCE loss. This indicates that our approach of capturing only differential information between AUs from each sample pair has higher efficiency for training encoder than methods that capture the entire facial information.
@@ 4.3.1 Compared with weighted cross-entropy loss
Model BL wi PS NS AU Index F1-Score(%) Accuracy(%)
1 2 4 6 7 10 12 14 15 17 23 24
A $\checkmark$ 39.8 44.3 58.8 76.0 77.7 76.9 88.0 64.5 49.0 59.0 43.5 43.4 60.1 79.1
B $\checkmark$ $\checkmark$ 40.1 43.0 58.9 75.4 79.3 78.8 87.8 64.1 48.7 61.0 46.9 47.5 61.00.9↑ 79.30.2↑
C $\checkmark$ $\checkmark$ $\checkmark$ 51.1 45.6 58.3 77.5 77.0 84.5 89.0 63.7 54.6 65.5 46.8 53.3 63.83.7↑ 79.80.7↑
D $\checkmark$ $\checkmark$ $\checkmark$ 50.2 43.0 58.7 77.9 77.3 83.7 88.7 64.2 53.9 64.9 50.0 56.4 64.14.0↑ 79.90.8↑
E $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 53.6 49.8 61.6 78.4 78.8 84.7 89.6 67.4 55.1 65.4 50.9 58.0 $\textbf{66.1}^{{\color[rgb]{1,0,0}6.0\uparrow}}$ $\textbf{80.4}^{{\color[rgb]{1,0,0}1.3\uparrow}}$
@@ 4.3.2 Evaluating other components
We further evaluate the effectiveness of weight coeffcient $\omega_{i}$, negative sample re-weighting strategy and positive sample samplying strategy. Quantitative results of our method with the above component combinations are summarized in Table 8. The baseline (BL) is rooted in SIMCLR [10], but positive sample pairs are changed to images with the highest similarity to original images. Other components are the weight coefficient ($w_{i}$) for both WCE and AUNCE, the positive sample sampling strategy (PS), and the negative sample re-weighting strategy (NS).
@@
Weight coefficient $w_{i}$: The weights $w_{i}$ have been pointed out as a fixed value in Section. 3.1, which achieve an increase of 0.9% in F1-score and 0.2% in Accuracy. For an AU, its activation indicator is a binary variable, and the corresponding information content can be measured using entropy: $-\log(p_{i})$. Therefore, AUs with lower activation probabilities carry more information and should be assigned higher weights. This forms the theoretical basis for our weighting strategy.
@@
Positive sample sampling strategy: As outlined in Table 8, the positive sample sampling strategy contributes to a 2.8% increase in F1-score and a 0.5% improvement in Accuracy compared to model B. This validates that selecting different types of positive sample pairs with appropriate probabilities can mitigate issues arising from AU labels tainted by noise and errors. The strategy achieves this by incorporating three types of positive sample pairs: the self-supervised signal pair, the highest similarity pair, and the mixed positive features pair. These pairs provide the models with diverse information from various views, thereby enhancing the generalization ability of the discriminative contrastive learning framework.
Negative sample re-weighting strategy: The negative sample re-weighting strategy to minority AUs results in a 3.1% increase in F1-score and a 0.6% increase in Accuracy compared to model B. This illustrates that the strategy can change the backpropagation gradient of minority and majority class samples, facilitating faster label category gradient updates with a smaller number of samples. Additionally, model E achieves an average increase of 2.3% in F1-score compared to model C, further supporting the efficacy of the re-weighting strategy.
@@
Figure 4: T-SNE visualization of feature representations on BP4D and DISFA datasets. Colors indicates whether AU1 exists. Top two rows: (a) Representations optimized by WCE on BP4D dataset, (b)-(f) Visualizations of ablation study on BP4D dataset, respectively represent model A-E. Bottom two rows: (g) Representations optimized by AUNCE on DISFA dataset, (h)-(l) Visualizations of ablation study on DISFA dataset, respectively represent model A-E.
@@
We further show the visualization using T-SNE dimensionality reduction of the learned feature representations optimized by AUNCE and WCE to qualitatively assess the efficacy of our proposed discriminative contrastive learning paradigm, as illustrated in Fig. 4. For each dataset, a random sample of 200 images is utilized for illustration purposes. A comparison between Fig. 4(a), Fig. 4(f) against Fig. 4(g), Fig. 4(l) reveals that the learned representations optimized by AUNCE exhibit greater separability than those optimized by WCE. Additionally, Fig. 4(b)-(f) and Fig. 4(g)-(l) qualitatively showcase the effectiveness of each component in our ablation study.
Based on all ablation studies, it demonstrates that our proposed discriminative contrastive learning paradigm enhances the encoding capability of the shared encoder, improves classification accuracy by addressing the class imbalance issue of each AU type, and bolsters model robustness against AU labels affected by noise and errors. This, in turn, leads to enhanced generalization performance in AU detection. The introduction of each component proves beneficial for performance improvement, with NS emerging as the most effective component. Moreover, the results on each AU are also shown in Table 8. It can be seen that the results on basically every AU have improved or approached after we add $w_{i}$, PS and NS.
@@ 4.4.1 $p_{1},p_{2},p_{3},p_{4}$, the probabilities of various positive sample pairs
The experimental results are presented in Table 9. When $p_{1}$=0.15, $p_{2}$=0.15, $p_{3}$=0.7, and $p_{4}$=0, the experimental result is optimal. It is observed that images with the lowest similarity to original images contribute minimally to the training process, as these images essentially represent noise and errors. The conclusion is consistent with [55]. The mixture of all positive samples plays a pivotal role, as it encourages the feature representation of each instance to be closer to the class centroid, facilitating the capture of essential characteristics of the positive class. The indispensability of the remaining two positive samples lies in their ability to mitigate the impact of noisy and false samples, thereby reinforcing the robustness of the training process.
@@ 5 Conclusion
In this paper, we proposes a discriminative contrastive learning framework tailored for AU detection, introducing a novel contrastive loss named AUNCE. AUNCE builds upon the InfoNCE loss, reformulating the task from learning pixel-level information of the entire face to focusing on the subtle feature variations that occur during AU activation, thereby enhancing detection efficiency. We further introduce a positive sample sampling strategy that constructs three types of positive sample pairs, effectively addressing the challenge of AU labels tainted by noise and errors in four AU datasets. To combat the class imbalance problem often seen in AU detection, we implement a negative sample re-weighting strategy specifically targeting minority AUs, facilitating the identification of hard negative samples. Experimental results reveal that AUNCE consistently outperforms existing methods, particularly those employing weighted cross-entropy loss. However, our negative sample reweighting strategy may inadvertently hinder the model’s ability to learn features of common AUs. An excessive emphasis on minority classes could result in underfitting of the majority classes. To overcome this limitation, our future research will focus on developing dynamic reweighting techniques with adaptive mechanisms that balance minority and majority classes, enhancing the robustness and adaptability of our framework across diverse conditions.
@@ Acknowledgement
This research was supported by the National Natural Science Foundation of China (Nos. 62176221, 61572407), and Sichuan Province Science and Technology Support Program (Nos. 2024NSFTD0036, 2024ZHCG0166). The computation is completed in the HPC Platform of Huazhong University of Science and Technology.

### v6

@@ 4.2.6 Cross dataset validation of AUNCE
WFL 63.81.4↑ 63.60.1↑
AUNCE $\textbf{64.1}^{{\color[rgb]{1,0,0}1.7\uparrow}}$ $\textbf{64.4}^{{\color[rgb]{1,0,0}0.9\uparrow}}$
MEGraph (Onestage) [26] 2022IJCAI WCE 63.6 60.8
WFL 64.20.6↑ 61.81.0↑
AUNCE $\textbf{66.1}^{{\color[rgb]{1,0,0}2.5\uparrow}}$ $\textbf{66.4}^{{\color[rgb]{1,0,0}5.6\uparrow}}$
@@ 4.3 Ablation Study
4.3.1 Compared with weighted cross-entropy loss and weighted focal loss
The primary innovation of this paper is the introduction of AUNCE, a novel contrastive loss designed to be compatible with various AU detection encoders. To assess its effectiveness, we compare the proposed AUNCE loss with traditional weighted cross-entropy loss (WCE) and the recently introduced weighted focal loss (WFL) [7] for AU detection during the pretraining phase, across multiple AU detection encoders. The results, presented in Table 8, demonstrate that AUNCE consistently outperforms the other losses in AU detection. Notably, the MEGraph encoder achieves the highest performance, with scores of 66.1% and 66.4% on BP4D and DISFA datasets, respectively, when trained using AUNCE. This highlights the advantage of emphasizing the capture of distinct features from individual samples, rather than solely relying on pixel-level information of the entire face in relation to AU detection.
Table 9: Comparison of the computational complexity and efficiency of the weighted cross-entropy loss (WCE), the weighted focal loss (WFL), and the proposed AUNCE loss on BP4D and DISFA dataset, where encoders of these three losses are both MEGraph with the first-stage training network.
Dataset Loss Complexity Frame Rate(FPS) 1 Epoch Time(s)
BP4D WFL $O(NC)$ 0.0293 2959
WCE $O(NC)$ 0.0271 2734
AUNCE $O(N+2)$ 0.0138 1388
DISFA WFL $O(NC)$ 0.0284 2473
WCE $O(NC)$ 0.0265 2308
AUNCE $O(N+2)$ 0.0136 1189
We compare the computational complexity and efficiency of the three losses. WCE and WFL both have an asymptotic complexity of $O(NC)$, as both of them involve computing probabilities for $C$ classes and summing over all $N$ samples, and WFL introducing negligible overhead due to the power and multiplication operations from the probability margin $\gamma$. In contrast, AUNCE has a lower complexity of $O(N+2)$, as it processes one anchor point, one positive sample, and $N$ negative samples, making it approximately twice as efficient as WCE and WFL when $C=2$. This suggests that AUNCE, by focusing on differential information between AUs, is more efficient for training the encoder compared to methods that capture entire facial information.
To further validate this advantage, we evaluate computational efficiency using two metrics: frame rate and training time per epoch. As shown in Table 9, under identical encoder parameters, AUNCE improves frame rate and reduces epoch training time by nearly half compared to WCE and WFL. These results confirm AUNCE’s superior efficiency in training, delivering comparable or better performance with significantly reduced computational overhead.
@@ 4.3.1 Compared with weighted cross-entropy loss and weighted focal loss
Model BL $\textbf{w}_{i}$ PS NS AU Index F1-Macro F1-Micro Accuracy
1 2 4 6 7 10 12 14 15 17 23 24 (%) (%) (%)
A $\checkmark$ 39.8 44.3 58.8 76.0 77.7 76.9 88.0 64.5 49.0 59.0 43.5 43.4 60.1 57.0 79.1
B $\checkmark$ $\checkmark$ 40.1 43.0 58.9 75.4 79.3 78.8 87.8 64.1 48.7 61.0 46.9 47.5 61.00.9↑ 57.20.2↑ 79.30.2↑
C $\checkmark$ $\checkmark$ $\checkmark$ 51.1 45.6 58.3 77.5 77.0 84.5 89.0 63.7 54.6 65.5 46.8 53.3 63.83.7↑ 57.60.6↑ 79.80.7↑
D $\checkmark$ $\checkmark$ $\checkmark$ 50.2 43.0 58.7 77.9 77.3 83.7 88.7 64.2 53.9 64.9 50.0 56.4 64.14.0↑ 57.70.7↑ 79.90.8↑
E $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 53.6 49.8 61.6 78.4 78.8 84.7 89.6 67.4 55.1 65.4 50.9 58.0 $\textbf{66.1}^{{\color[rgb]{1,0,0}6.0\uparrow}}$ $\textbf{58.2}^{{\color[rgb]{1,0,0}1.2\uparrow}}$ $\textbf{80.4}^{{\color[rgb]{1,0,0}1.3\uparrow}}$
@@ 4.3.2 Evaluating other components
We further assess the effectiveness of all proposed components. Quantitative results of ablation study are summarized in Table 10. The baseline (BL) is based on SIMCLR [13], with positive sample pairs replaced by images most similar to the original ones. Other components include the weight coefficient ($w_{i}$) for both WCE and AUNCE, the positive sample sampling strategy (PS), and the negative sample re-weighting strategy (NS).
@@
Weight coefficient $w_{i}$: The weights $w_{i}$ have been pointed out as a fixed value in Section. 3.1, which achieve an increase of 0.9% in F1-macro score, 0.2% in F1-micro score and 0.2% in Accuracy. For an AU, its activation indicator is a binary variable, and the corresponding information content can be measured using entropy: $-\log(p_{i})$. Therefore, AUs with lower activation probabilities carry more information and should be assigned higher weights. This forms the theoretical basis for our weighting strategy.
@@
Positive sample sampling strategy: As shown in Table 10, the positive sample sampling strategy improves the F1-macro score by 2.8%, the F1-micro score by 0.4%, and Accuracy by 0.5% compared to model B. This demonstrates that selecting diverse positive sample pairs with appropriate probabilities helps mitigate issues related to noisy or false AU labels. The strategy utilizes three types of positive sample pairs: self-supervised signal pairs, highest similarity pairs, and mixed feature pairs. By incorporating these varied pairs, the model benefits from more informative inputs, enhancing its generalization ability and discriminative power. These improvements enable more effective AU detection, even in the presence of label noise and class imbalance, reinforcing the model’s robustness across various AU detection scenarios.
(a) Visualizing feature maps generated by model E (With re-weighting strategy)
(b) Visualizing feature maps generated by model C (Without re-weighting strategy)
Figure 4: Visualizing feature maps generated by models with and without the negative sample re-weighting strategy on BP4D dataset, which can help illustrate how the model with the negative sample re-weighting strategy concentrates more on the related facial regions. (a): Visualizing feature maps generated by model E (With the negative sample re-weighting strategy), and (b): Visualizing feature maps generated by model C (Without the negative sample re-weighting strategy).
Negative sample re-weighting strategy: The negative sample re-weighting strategy, which prioritizes minority AUs during training, improves the F1-macro score by 3.1%, the F1-micro score by 0.5%, and accuracy by 0.6% compared to model B, demonstrating its ability to adjust backpropagation gradients for minority and majority classes, enabling faster updates with fewer samples. Additionally, model E achieves a 2.3% average increase in F1-macro score compared to model C, further validating the strategy’s effectiveness. As shown in Fig. 4, model E (top row) focuses more on facial regions relevant to rare AUs, enhancing the detection of these challenging cases and reinforcing its utility in addressing class imbalance.
@@
Figure 5: T-SNE visualization of feature representations on BP4D and DISFA datasets. Colors indicates whether AU1 exists. Top row: (a) Representations optimized by WCE on BP4D dataset, (b)-(f) Visualizations of ablation study on BP4D dataset, respectively represent model A-E. Bottom row: (g) Representations optimized by AUNCE on DISFA dataset, (h)-(l) Visualizations of ablation study on DISFA dataset, respectively represent model A-E.
@@
To qualitatively assess the efficacy of our discriminative contrastive learning paradigm, we visualize the learned feature representations optimized by AUNCE and WCE using T-SNE dimensionality reduction, as shown in Fig.5. A random sample of 200 images from each dataset is used for illustration. Comparing Fig.5(a), Fig.5(f) with Fig.5(g), Fig.5(l) demonstrates that features learned with AUNCE show greater separability than those optimized by WCE. Additionally, Fig.5(b)-(f) and Fig. 5(g)-(l) highlight the contribution of each component in our ablation study.
The ablation studies confirm that the proposed discriminative contrastive learning paradigm enhances the encoder’s feature representation capability, addresses class imbalance for each AU type, and improves robustness to noisy and false labels, ultimately boosting generalization performance in AU detection. Each component contributes to performance gains, with NS proving most effective. Furthermore, as shown in Table 10, the addition of $w_{i}$, PS, and NS consistently improves or maintains performance across nearly all AUs.
The experimental results are presented in Table 11. When $p_{1}$=0.15, $p_{2}$=0.15, $p_{3}$=0.7, and $p_{4}$=0, the experimental result is optimal. It is observed that images with the lowest similarity to original images contribute minimally to the training process, as these images essentially represent noise and errors. The mixture of all positive samples plays a pivotal role, as it encourages the feature representation of each instance to be closer to the class centroid, facilitating the capture of essential characteristics of the positive class. The indispensability of the remaining two positive samples lies in their ability to mitigate the impact of noisy and false samples, thereby reinforcing the robustness of the training process.
However, hyperparameter tuning is a limitation of our paper. Although the proposed method yields promising results, we observed that the performance is highly sensitive to the values of $p_{1}$, $p_{2}$, and $p_{3}$, with their optimal settings varying across different datasets. Moreover, the parameter tuning process can only be guided by experience gained from extensive experimentation, which results in a tuning scheme that lacks flexibility.
@@ 5 Conclusion
In this paper, we propose a discriminative contrastive learning framework tailored for AU detection, introducing a novel contrastive loss named AUNCE. Building upon the InfoNCE loss, AUNCE refines the task by shifting focus from learning pixel-level information of the entire face to capturing subtle feature variations during AU activation, thus improving detection efficiency. We also introduce a positive sample sampling strategy that constructs three types of positive sample pairs, selectively augmenting positive samples, especially in the presence of noisy and false AU labels. This approach ensures the model is trained on more reliable examples, improving upon conventional methods that lack adaptive sampling. To address class imbalance in AU detection, we implement a negative sample re-weighting strategy that prioritizes minority AUs, helping the model focus on hard-to-detect samples. This strategy promotes hard example mining by giving more weight to underrepresented classes, enhancing model performance and generalization. Experimental results demonstrate that AUNCE consistently outperforms existing methods, particularly those utilizing weighted cross-entropy loss.
However, hyperparameter tuning remains a limitation of our study. While the proposed method demonstrates promising results, we found that performance is highly sensitive to the values of certain hyperparameters, with optimal settings varying across different datasets. The parameter tuning process is largely driven by empirical experience gained from extensive experimentation, which results in a tuning scheme that lacks the necessary flexibility. To address this limitation, our future work will focus on the development of adaptive hyperparameter selection techniques that can dynamically adjust to different datasets and conditions, enabling more robust and generalized performance across a variety of tasks and settings.
@@ Acknowledgement
This research was supported by the National Natural Science Foundation of China (Nos. 62176221, 61572407), and Sichuan Province Science and Technology Support Program (Nos. 2024NSFTD0036, 2024ZHCG0166).